40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 7 20:43:35 2009, Clara, Jenne, Rana, Jan, Update, PEM, Two Guralps plugged in, prepped for huddle test Untitled.png
    Reply  Sat Aug 8 17:19:07 2009, rana, Update, PEM, Two Guralps plugged in, prepped for huddle test Untitled.png
Message ID: 1868     Entry time: Sat Aug 8 17:19:07 2009     In reply to: 1866
Author: rana 
Type: Update 
Category: PEM 
Subject: Two Guralps plugged in, prepped for huddle test 

 I found that several of the cables are unlabeled so I'm not sure what's plugged in. In the end, I found that the TEMP_2, _3, & _44 channels were working and so I plugged in anything that looked seismic into there.

TEMP_2 is now apparently the X channel of the 2nd Guralp. If someone can figure out which cable belongs to the Y channel, please plug it into TEMP_3 and then we can fix the channel names.

I also removed (gently) all of the accelerometers from MC2's chamber. This didn't break the lock, but I intentionally broke it to make sure it reacquired fine. It did and the MC TRANS QPD showed no significant shift afterwards.

Attachment 1: Untitled.png  80 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-